jbell:

Solve the equation. –9x + 1 = –x + 17 x = –8 x = –2 x = 2 x = 8

xxDeppressionxx:

Isolate the variable by dividing each side by factors that don't contain the variable. x=−2

umm:

Welcome to QuestionCove, jbell. Given\[−9x+1=−x+17\]Add x to both sides\[−9x+1+x=−x+17+x\]\[−8x+1=17\]Subtract both sides by 1\[−8x+1−1=17−1\]\[−8x=16\]Divide -8 from both sides\[\frac{ -8x }{ -8 }=\frac{ 16 }{ 8 }\]\[x=−2\]Therefore \[x=−2\]

sxdsouls:

Step 1: Add x to both sides. −9x+1+x=−x+17+x −8x+1=17 Step 2: Subtract 1 from both sides. −8x+1−1=17−1 −8x=16 Step 3: Divide both sides by -8. −8x −8 = 16 −8 x=−2 Sorry if this is Direct !! tried to help

jbell:

ok I'm not on here for long

sxdsouls:

oop lmao

sxdsouls:

but welcome,

xxDeppressionxx:

Welcome, jbell to Qc